OpenStudy (anonymous):

A bus travels from Oakridge to Hillsboro, a distance of 320 miles. The bus is traveling at 60 mph. The driver is required to contact the station in Hillsboro when he is between 50 and 100 miles from the station. During what time interval should he contact the station at Hillsboro.

OpenStudy (anonymous):

60 miles per hour 1 mile per minute 320 minutes for the trip 50 minutes to 100 minutes before the end of the trip 320 - 50 = 270 320 - 100 = 220 220 min and 270 min or 3 h 40 min and 4 h 30 min

OpenStudy (anonymous):

50 to 100 miles from city means from 220 mile to 270 mile at 60mph bus will take t=distance/ speed=220/60 to t=270/60

OpenStudy (anonymous):

thanks so much